OpenStudy (anonymous):

Finding the two-sided limit of the following general form: \[ \lim_{x→0} (1 − ax)^{1/x}\] What's the trick here? A medal will be given for showing the steps clearly :-)

OpenStudy (anonymous):

\[\large y=(1-ax)^{\frac{1}{x}} \] \[\large lny=\frac{1}{x}ln(1-ax) \] now take the limit using L'Hopital's rule.

OpenStudy (fwizbang):

The trick here is to write \[ (1-ax)^{1/x}=e ^{ \ln(1-ax)/x}\] Apply l'hopital's rule to the exponent as x goes to zero ln(1-ax)/ x ---> -a/(1-ax) ----> -a so that \[\[ (1-ax)^{1/x}=e ^{ -a}\]

OpenStudy (anonymous):

Gj both of you and ty :-)